OpenStudy (testing123guest):

BRAIN TEZER OF THE DAY Shaun is 4 years older than her sister Charmaine. If the sum of their ages is 16 how old is Charmaine?

OpenStudy (anonymous):

Four years old!!!

OpenStudy (calculusxy):

shaun = 4 + c charmaine = c c + (4 +c) = 16 2c + 4 = 16 2c = 12 c = 6 check: charmaine = 6 shaun = 6 + 4 = 10 10 + 6 = 16 <- correct

OpenStudy (testing123guest):

16 is not one of the answers @love.melly was..........................................................CORRECT!!!!!!!!!!!

OpenStudy (calculusxy):

i never said 16 is one of the answers. i just checked to see whether my answer was correct or not. i said that charmaine was 6 years old.

OpenStudy (testing123guest):

OH sorry u are correct then sorry man lol

OpenStudy (calculusxy):

but if charmaine was 4, then shaun would be 10 (6 + 4). if you add them up you get 14, not 16.

OpenStudy (testing123guest):

so is it 4 , 6 , or 10

OpenStudy (calculusxy):

i said that \(\text{Charmaine = 6 years old}\)